# Accelerating AI Performance using Anderson Extrapolation on GPUs

### v1

@@ Abstract
We present a novel approach for accelerating AI performance by leveraging Anderson extrapolation, a vector-to-vector mapping technique based on a window of historical iterations. By identifying the crossover point where a mixing penalty is incurred, the method focuses on reducing iterations to convergence, with fewer more compute-intensive but generally cacheable iterations, balancing speed and memory usage with accuracy and algorithmic stability, respectively. We demonstrate significant improvements, in both training and inference, motivated by scalability and efficiency extensions to the realm of high-performance computing (HPC).
@@ 1 Introduction
Anderson extrapolation [1, 2, 27, 33, 16] has recently been applied to deep equilibrium models (DEQs) [7, 8, 9, 10, 24, 17]. Kolter et al. [34] found the gains not substantial due to early termination with a loose convergence tolerance. They focused on Anderson extrapolation during training. Here, we show significant acceleration of AI performance with Anderson on GPUs for both the forward pass (running inferences faster) and training (generating models faster). We demonstrate acceleration of the forward pass with standard Anderson as a baseline for future work with stochastic variants [30] and accelerating the backward pass with Jacobian-free methods like Jacobian-Free Backpropagation (JFB) and Neumann series gradient approximations [16].
@@
As AI demand grows, as shown in Fig. 2 [3, 15, 28, 25], high-performance computing (HPC) is becoming critical due to economic pressures from the growth of data and AI infrastructure [29]. Low-memory acceleration techniques, like Anderson extrapolation, will be key to increasing HPC-based AI computational efficiency. This study investigates matrix-free Anderson extrapolation on GPUs, emphasizing gains from advanced computing architectures compared to CPUs. Our goal is to maximize computational efficiency while reducing iterations to convergence by reusing previous iterations to avoid unnecessary gradient calculations, gaining benefits expected from second-order methods (e.g., [32]) without manipulating Hessian matrices.
@@ 1.1 Leveraging extrapolation for AI and HPC advances
Anderson extrapolation, a windowing technique for accelerating nonlinear fixed point iterations, is widely applied in fields like density functional theory, kinetic theory, and climate spin-up. It is well-suited for distributed memory parallelization and GPU implementation. It is a staple of major open-source large-scale solver libraries, including PETSc [11, 12], SUNDIALS [23], Trilinos [19, 22, 21, 20], and deal.II [13, 4, 5, 6]. It can be applied to machine learning training, smoothing out standard forward iterations and achieving superior accuracy in training and testing error. Benchmarking results on CIFAR10 show expected robustness benefits and allow characterization of the temporal advantages or disadvantages from the higher cost per iteration, where a small residual minimization step is applied at each new function evaluation.
@@
Figure 4: Deep equilibrium neural network model architecture (Source: NeurIPS Tutorial, 2020 [34]). $f(z,x)=\mathrm{norm}(\mathrm{ReLU}(z+\mathrm{norm}(x+W_{2}*(\mathrm{norm}(\mathrm{ReLU}(W_{1}*z))))))$norm" here is a group norm, representing a statistical normalization [31].
1.2 Balancing memory and stability
@@ 1.3 Deep equilibrium neural network models
Deep equilibrium models (DEQs) are the continuum limit of explicit neural networks as the number of layers approaches infinity [26], approximating many explicit layers with a single, implicit layer with exponentially fewer parameters using a backward pass including the output. This reduces the inverse problem in parameter space to a fixed-point iteration problem, enabling the usage of nonlinear, vector-to-vector mapping techniques to compute the fixed-point iterations that converge to the deep equilibrium state parameters by minimizing the loss function. With gains in memory and acceleration, DEQs are fit for large-scale computer vision and natural language processing tasks and benefit more from matrix-vector operation-optimized computing architectures like GPUs and CPU-GPU superchips.
The standard approach using forward iteration for fixed-point iteration problems often does not efficiently converge to the fixed point and suffers from initially slow error reduction and local minimum trapping in nonlinear problems like deep neural networks. Vector-to-vector mapping techniques like Anderson extrapolation outperform standard forward iteration by combining information from previous iterations to span a searchable subspace to extrapolate the next iteration, enhancing convergence rates at the expense of memory usage in each iteration.
@@
DEQs provide a framework for accelerating deep learning, extending the capacity of deep networks within a single-layer architecture through fixed-point computations and advanced root-finding algorithms. Their amenability to convergence acceleration with techniques like Anderson positions DEQs as a robust method to reduce computation needed to build state-of-the-art models and scale up beyond current computational limitations.
@@ 2 Methods
This work demonstrates Anderson extrapolation to accelerate AI performance algorithmically without increasing processors. Since it is a vector-to-vector mapping technique, it benefits from hardware optimized for uniform operations, like GPUs. We benchmark Anderson acceleration against standard forward iteration on GPUs and CPUs.
@@ 2.1 Mathematical formulation
Fixed-point acceleration starts with the fixed point iteration formula $z^{\star}=f(z^{\star},x)$. Forward iteration, $z^{k+1}=f(z^{k},x)$, moves step-wise towards this fixed point.
Anderson acceleration uses a linear combination of prior iterates, $z^{k+1}=\sum_{i=1}^{m}\alpha_{i}f(z^{k-i+1},x)$, optimizing $\alpha_{i}$ to minimize the residual norm, $\frac{\|f(z^{k},x)-z^{k}\|_{2}}{\|f(z^{k},x)\|_{2}+\lambda}$, leading to faster convergence. The coefficients must sum to unity:
@@
Anderson acceleration can incorporate a mixing parameter $\beta$:
@@ 2.4 GPU Optimization and Parallelization
GPUs, suited for uniform tasks with high throughput, map well with Anderson acceleration. This work combines suitable algorithms with appropriate architecture to enhance performance without upgrading hardware or using more processors.
@@ 3 Results
This work demonstrates that Anderson extrapolation has a higher cost per iteration, measured in function evaluations or epochs. The main benefit is that Anderson extrapolation exhibits less fluctuation in accuracy, as seen in the test accuracy, whereas forward iteration shows more significant ups and downs in both training and testing accuracy, potentially indicating overfitting. Anderson acceleration reaches a higher accuracy plateau for both training and test datasets, suggesting better generalization capability.
@@
Implicit neural network model architecture performance is analyzed with different vector-to-vector mapping techniques. The goal is to understand how incorporating Anderson acceleration impacts model accuracy and performance. The stability of train and test accuracy is observed, and Anderson acceleration demonstrates higher consistency over numerous epochs, whereas forward iteration reveals significant swings in train and test accuracy. Initialization error with Anderson is lower than with forward iteration.
Anderson acceleration reaches higher accuracies in training and testing in less time than forward iteration. Anderson acceleration is also superior with random inputs. Across testing with random inputs, Anderson acceleration consistently outperforms or matches forward iteration, depending on target relative residual accuracy.
@@ 4 Discussion
GPUs have been shown to accelerate Anderson extrapolation beyond what could be achieved with standard forward iterations or with Anderson on CPUs. This is notable before reaching the ‘crossover point,’ the trade-off between computation speed and accuracy, illustrated in Fig. 1 and Fig. 6. The ‘mixing penalty’ due to the additional computational cost associated with Anderson acceleration is offset by the parallel processing capabilities of GPUs, enabling faster convergence than with CPUs or standard forward iterations alone.
Figure 6: Evaluating relative residual, $\frac{\|f(z^{k},x)-z^{k}\|_{2}}{\|f(z^{k},x)\|_{2}+\lambda}$, for a random input $x$. A typical GPU is $\sim$100-150x faster to target relative residual than a typical CPU using Anderson, with a mixing penalty that is $\sim 10^{-1}-10^{-2}$ lower.
@@
Figure 7: Deep equilibrium model is $\sim$10x faster to stable convergence with Anderson relative to standard forward iteration, per Tab. 1
@@ 5 Conclusion
The integration of Anderson acceleration within deep learning workflows presents substantial improvements in computational efficiency, accuracy, and generalizability of implicit neural networks. Porting and parallelizing vector-to-vector mapping techniques onto emerging CPU-GPU hybrid architectures holds promise. The accuracy and speed of deep equilibrium neural network training and inferences could be improved further, making them more viable for real-world applications beyond the classification task demonstrated herein. Based on investigations of explicit and implicit memory requirements [26], optimizations based on an Anderson-accelerated, fixed-point iteration implicit memory approach [34] are effective in memory-intensive computer vision processing, reducing memory and bandwidth consumption without compromising performance [26].
@@ 6 NeurIPS Limitation and Broader Impact Statements
SAA acknowledges funding from the KAUST Fellowship, the ECRC, and Prof. David Keyes throughout the duration of this work. SAA wishes to express sincere gratitude to Henning Soller, Mohamed-Slim Alouni, Matteo Parsani, George Turkiyyah, and Frederic Laquai for their invaluable chats throughout this research. Special gratitude is extended to my family—late grandparents, parents, wife, daughter, and brothers—and to my friends for their unwavering support and encouragement throughout the duration of this study.

### v2

@@ Abstract
We present a novel approach for accelerating AI performance by leveraging Anderson extrapolation, a vector-to-vector mapping technique based on a window of historical iterations. By identifying the crossover point (Fig. 1) where a mixing penalty is incurred, the method focuses on reducing iterations to convergence, with fewer more compute-intensive but generally cacheable iterations, balancing speed and memory usage with accuracy and algorithmic stability, respectively. We demonstrate significant improvements in both training and inference, motivated by scalability and efficiency extensions to the realm of high-performance computing (HPC).
@@ 1 Introduction
Anderson extrapolation [1, 2, 27, 30, 34, 16] has recently been applied to deep equilibrium models (DEQs) [7, 9, 8, 10, 24, 17]. Kolter et al. [35] found the gains not substantial due to early termination with a loose convergence tolerance. They focused on Anderson extrapolation during training. Here, we show significant acceleration of AI performance with Anderson on GPUs for both the forward pass (running inferences faster) and training (generating models faster). We demonstrate acceleration of the forward pass with standard Anderson as a baseline for future work with stochastic variants [31] and accelerating the backward pass with Jacobian-free methods like Jacobian-Free Backpropagation (JFB) and Neumann series gradient approximations [16].
@@
As AI demand grows, as shown in Fig. 2 [3, 15, 28, 25], high-performance computing (HPC) is becoming critical due to economic pressures from the growth of data and AI infrastructure [29]. Low-memory acceleration techniques, like Anderson extrapolation, will be key to increasing HPC-based AI computational efficiency. This study investigates matrix-free Anderson extrapolation on GPUs, emphasizing gains from advanced computing architectures compared to CPUs. Our goal is to maximize computational efficiency while reducing iterations to convergence by reusing previous iterations to avoid unnecessary gradient calculations, gaining partial benefits expected from second-order methods (e.g., [33]) without manipulating Hessian matrices.
@@ 1.1 Leveraging extrapolation for AI and HPC advances
Anderson extrapolation, a windowing technique for accelerating nonlinear fixed point iterations diagrammed in Figs. 3 and 4, is widely applied in fields like density functional theory, kinetic theory, and climate spin-up. It is well-suited for distributed memory parallelization and GPU implementation. It is a staple of major open-source large-scale solver libraries, including PETSc [11, 12], SUNDIALS [23], Trilinos [19, 22, 21, 20], and deal.II [13, 4, 5, 6]. It can be applied to machine learning training, smoothing out standard forward iterations and achieving superior accuracy in training and testing error. Benchmarking results on CIFAR10 show expected robustness benefits and allow characterization of the temporal advantages or disadvantages from the higher cost per iteration, where a small residual minimization step is applied at each new function evaluation.
@@
Figure 4: Deep equilibrium neural network model architecture (Source: NeurIPS Tutorial, 2020 [35]). $f(z,x)=\mathrm{norm}(\mathrm{ReLU}(z+\mathrm{norm}(x+W_{2}*(\mathrm{norm}(\mathrm{ReLU}(W_{1}*z))))))$. "norm" here is a group norm, representing a statistical normalization [32].
1.2 Balancing memory and convergence rate
@@ 1.3 Deep equilibrium neural network models
Deep equilibrium models (DEQs) are the continuum limit of explicit neural networks as the number of layers approaches infinity [26], approximating many explicit layers with a single, implicit layer with exponentially fewer parameters using a backward pass including the output. This reduces the inverse problem in parameter space to a fixed point iteration problem, enabling the usage of nonlinear, vector-to-vector mapping techniques to compute the fixed point iterations that converge to the deep equilibrium state parameters by minimizing the loss function. With gains in memory and acceleration, DEQs are fit for large-scale computer vision and natural language processing tasks and benefit more from matrix-vector operation-optimized computing architectures like GPUs and CPU-GPU superchips.
The standard approach using forward iteration for fixed point iteration problems often does not efficiently converge to the fixed point and suffers from initially slow error reduction and local minimum trapping in nonlinear problems like deep neural networks. Anderson extrapolation outperforms standard forward iteration by combining information from previous iterations to span a searchable subspace to extrapolate the next iteration, enhancing convergence rates at the expense of memory usage in each iteration. When the original fixed point iteration is contractive and thus guaranteed to converge, Anderson is theoretically guaranteed not to be slower [30] and it experimentally observed to be considerably faster in numerous applications.
@@
DEQs provide a framework for accelerating deep learning, extending the capacity of deep networks within a single-layer architecture through fixed point computations and advanced root-finding algorithms. Their amenability to convergence acceleration with techniques like Anderson positions DEQs as a robust method to reduce computation needed to build state-of-the-art models and scale up beyond current computational limitations.
@@ 2 Methods
This work demonstrates Anderson extrapolation to accelerate AI performance algorithmically without increasing processors. Since it does not require inverting matrices approximating Hessians of the dimension of the state space, but only matrices of the dimension of the Anderson window size, it benefits from hardware optimized for uniform vector operations, like GPUs. We benchmark Anderson acceleration against standard forward iteration on GPUs and CPUs.
@@ 2.1 Mathematical formulation
Fixed point acceleration starts with the fixed point iteration formula $z^{\star}=f(z^{\star},x)$. Forward iteration, $z^{k+1}=f(z^{k},x)$, moves step-wise towards this fixed point.
Anderson acceleration uses a linear combination of prior iterates, $z^{k+1}=\sum_{i=1}^{m}\alpha_{i}f(z^{k-i+1},x)$, optimizing $\alpha_{i}$ to minimize the residual norm, $\frac{\|f(z^{k},x)-z^{k}\|_{2}}{\|f(z^{k},x)\|_{2}+\lambda}$, leading to faster convergence. The coefficients must sum to unity, thus:
@@
Anderson acceleration generally includes a mixing parameter $\beta$, incorporating some inertia when $\beta<1$:
@@ 2.4 GPU Optimization and Parallelization
Anderson acceleration maps well to GPUs, suited for uniform tasks with high throughput.
@@ 3 Results
Anderson extrapolation has a higher cost per iteration, measured in function evaluations or epochs. The main benefit is that Anderson extrapolation exhibits less fluctuation in accuracy, as seen in the test accuracy, whereas forward iteration shows more significant ups and downs in both training and testing accuracy, potentially indicating overfitting. Anderson acceleration reaches a higher accuracy plateau for both training and test datasets, suggesting better generalization capability. We speculate that this is due to better avoidance of suboptimal local minima because of the wider window of trial vectors from which each step is drawn.
@@
Implicit neural network model architecture performance is analyzed with the goal of understanding how incorporating Anderson acceleration impacts model accuracy and performance. The stability of train and test accuracy is observed, and Anderson acceleration demonstrates higher consistency over numerous epochs, whereas forward iteration reveals significant swings in train and test accuracy. Initialization error with Anderson is lower than with forward iteration.
We observe Anderson acceleration reaches higher accuracies in training and testing in less time than forward iteration. Anderson acceleration is also superior with random inputs. Across testing with random inputs, Anderson acceleration consistently outperforms or matches forward iteration, depending on target relative residual accuracy.
@@ 4 Discussion
GPUs have been shown to accelerate Anderson extrapolation beyond what could be achieved with standard forward iterations or with Anderson on CPUs. This is notable before reaching the ‘crossover point,’ the trade-off between computation speed and accuracy, illustrated in Figs. 1 and 6. The ‘mixing penalty’ due to the additional computational cost associated with Anderson acceleration is offset by the parallel processing capabilities of GPUs, enabling faster convergence than with CPUs or standard forward iterations alone.
Figure 6: Evaluating relative residual, $\frac{\|f(z^{k},x)-z^{k}\|_{2}}{\|f(z^{k},x)\|_{2}+\lambda}$, for a random input $x$. A typical GPU is approximately 100-150x faster to target relative residual than a typical CPU using Anderson, with a mixing penalty that is approximately $10^{-1}$ to $10^{-2}$ lower.
@@
Figure 7: Deep equilibrium model is approximately 10x faster to stable convergence with Anderson relative to standard forward iteration, per Table 1
@@ 5 Conclusion
The integration of Anderson acceleration within deep learning workflows presents substantial improvements in computational efficiency, accuracy, and generalizability of implicit neural networks. Porting and parallelizing matrix-free acceleration techniques onto emerging CPU-GPU hybrid architectures holds promise. The accuracy and speed of deep equilibrium neural network training and inferences could be improved further, making them more viable for real-world applications beyond the classification task demonstrated herein. Based on investigations of explicit and implicit memory requirements [26], optimizations based on an Anderson-accelerated, fixed-point iteration implicit memory approach [35] are effective in memory-intensive computer vision processing, reducing memory and bandwidth consumption without compromising performance [26].
@@ 6 NeurIPS Limitation and Broader Impact Statements
SAA acknowledges funding from the KAUST Fellowship and the Extreme Computing Research Center throughout the duration of this work. SAA wishes to express sincere gratitude to Henning Soller, Mohamed-Slim Alouni, Matteo Parsani, George Turkiyyah, and Frédéric Laquai for their invaluable chats throughout this research. Special gratitude is extended to my family—late grandparents, parents, wife, daughter, and brothers—and to my friends for their unwavering support and encouragement throughout the duration of this study.